A charge of +3.0 𝑥 10−5𝐶 is 10cm from a charge of +5.0 𝑥10−6𝐶. Calculate 

the magnitude of the repulsive force between the charges.


1
Expert's answer
2021-02-10T10:06:57-0500
"F=\dfrac{kq_1q_2}{r^2},""F=\dfrac{9\cdot10^9\ \dfrac{Nm^2}{C^2}\cdot3\cdot10^{-5}\ C\cdot5\cdot10^{-6}\ C}{(0.1\ m)^2}=135\ N."